# PII-Scope: A Benchmark for Training Data PII Leakage Assessment in LLMs

### v1

@@ 6.2 Continual PII Extraction
Figure 11: Impact of using synthetic data as the adversary’s knowledge in PII attacks. We use synthetic data at varying levels (pruple and green bars) in place of real data (yellow bars) from $\mathcal{D}_{adv}$. For the ICL attack Huang et al. (2022a), we fix the number of demonstrations at 4 and run the demonstration selection process using 21 different seeds from a pool of 64 synthetic examples. In the PII Compass attack Nakka et al. (2024), we set the prefix length to 50 tokens and iterate over 64 synthetic prefixes (see Figures 27 and  28). For the SPT attack Kim et al. (2024), we repeat the experiment with 20 task-aware prompt initializations, as shown in Figure 19 in the Appendix.
@@ 7 Ablation Studies
For example, given a real PII pair {Karen Arnold, klarnold@flash.net} in the adversary dataset $\mathcal{D}_{\text{adv}}$ as shown in Figures 21 and 22, we generate synthetic PII pairs in two variations: 1. Altering only the name with email-domain retained (e.g., {"Cameron Thomas", "cthomas@flash.net"}, as shown in Figures 23 and 24 in the Appendix). 2. Altering both the name and the domain with synthetic ones (e.g., {"Cameron Thomas", "cthomas@medresearchinst.org"}, as shown in Figures 25 and 26 in the Appendix).
For synthetic prefixes in the PII-Compass attack (Nakka et al., 2024), we use GPT-3.5 (OpenAI, 2023) to generate email conversation sentences of 50 tokens in length between employees of an energy corporation like Enron, as illustrated in Figures 27 and 28.
The results of PII attacks on these synthetic data experiments are presented in Figures 11 for ICL, PII-Compass, and SPT attacks in three columns, respectively. Overall, our observations are as follows: 1. When both the name and domain are replaced with synthetic data, the extraction rates for both ICL and SPT attacks are notably lower (shown in purple bars) compared to the original performance with real PII pairs (shown in yellow bars). 2. When only the name part is anonymized, the performance of the ICL attack (shown in green bars) remains closer to the original performance with real PII pairs (shown in yellow bars). In contrast, the performance of SPT attacks in this setting shows a significant drop in performance (shown in green bars) from that with original PII pairs (shown in yellow bars) and in fact, the SPT attack, does not even surpass the performance of simple template prompting, as shown in Figure 7(a). 3. With synthetic prefixes generated by GPT (OpenAI, 2023), the performance (shown in purple bars) is substantially lower than the original performance with real prefixes from subjects in $\mathcal{D}_{adv}$, as illustrated in Figure 11(c). Our experiments suggests that for effective PII extraction with PII-Compass, having a prefix that closely resembles the true domain is essential.
Figure 12: Impact of order of subjects in the demonstration prompt of the ICL attack. We first select $k=\{2,4,6,8,16,32,64\}$ PII pairs from the pool of $M=64$ PII pairs in $\mathcal{D}_{adv}$ using a single seed. Next, we vary the order of the $k$ demonstrations by generating 20 different permutations for each $k$. We visualize the box plot of extraction rates across these 20 different permutations and observe that the ICL attack (Huang et al., 2022a) shows increased sensitivity to demonstration order as the number of demonstrations $k$ increases.
@@
Figure 13: Soft-prompt transferability. The Y-axis denotes the template structure used for training the soft prompt embeddings. The X-axis shows the four target templates used during the attack stage. To conduct this study, we prepend the trained soft prompt embeddings from different source templates (indicated along the Y-axis) to different target template prompts (indicated along the X-axis) and report the average PII extraction performance over 21 soft-prompt initializations shown in Figure 19.
@@
(b) PII Compass attack
Figure 14: True-prefix attack and Template attack on the finetuned model. On the left, we show the performance of the true-prefix attack (Carlini et al., 2021a), and on the right, we present the performance of the template attack (Huang et al., 2022a). Results for the pretrained model are shown in light color, while results for the finetuned model are shown in dark color. Across the board, we observe that PII extraction rates on the finetuned model are significantly higher than those on the pretrained model.
(a) ICL attack
(b) PII Compass attack
Figure 15: Sensitivity of Hard-Prompt Attacks on the Finetuned Model. Similar to the results on the pretrained model in Figure 7, the ICL attack (Huang et al., 2022a) on the left shows sensitivity to the selection of demonstrations from the available pool of $\mathcal{D}_{adv}$, while the PII Compass attack (Nakka et al., 2024) on the right illustrates the impact of varying true prefixes from other data subjects in $\mathcal{D}_{adv}$.
@@ 8 PII Attacks on Finetuned Model
Figure 16: Sensitivity of SPT Attack (Kim et al., 2024) on the Finetuned Model. We examine the variation in PII extraction rates by analyzing the impact of three independent factors. Each factor is varied independently from the base configuration, and the results show that the SPT attack requires careful hyperparameter selection for optimal performance.
@@
Continual PII extraction. We also conduct continual PII extraction on the finetuned model by leveraging successfully extracted PII pairs along with the originally available PII pairs in $\mathcal{D}_{adv}$. We perform this experiment with 5 task-aware initializations (see first 5 in Figure 19 in the Appendix) for each template. From results in Figure 17, we observe that the average extraction rates improve for templates A, B, C, and D from 9.09%, 19.9%, 24.1%, 28.2% at the end of round 1 to 12.1%, 35.8%, 39.5%, 42.1% at the end of round 2. All templates achieve a boost of more than 1.5x, except for template A, which shows greater variance in extraction rates across different initializations.
@@ 4 Templates: [A, B, C, D])
Table 5: Evaluating PII attacks with higher query budgets on the finetuned model. The first four columns outline the threat setting in terms of data access in $\mathcal{D}_{adv}$ and $\mathcal{D}_{eval}$. The fifth column shows the model access type (W.B.: white box, B.B.: black box). We conduct PII attacks by querying the model multiple times, either through simple top-$k$ model sampling or by varying configuration settings within each attack method. Unlike attacks on the pretrained model, even the simple template attack (Huang et al., 2022a) achieves more than 50% accuracy in finetuned settings. Furthermore, similar to earlier results on the pretrained model, we observe that the extraction rate improves by 1.2x-2.8x compared to the best extraction rate observed with a single query.
@@
Figure 17: Continual PII extraction on the finetuned model. We report the extraction rates of the SPT attack (Kim et al., 2024) over ten rounds for four templates in a continual learning setting. At the end of each round, successfully extracted PIIs are incorporated to retrain the soft prompt embeddings for the subsequent round. The average extraction rate, along with its range, is plotted for the first five soft-prompt initializations shown in Figure 19.

### v2

@@ 7.2 Continual PII Extraction
Figure 11: Impact of using synthetic data as the adversary’s knowledge in Email PII attacks on pretrained GPTJ-6B. We use synthetic data at varying levels (purple and green bars) in place of real data (yellow bars) from $\mathcal{D}_{adv}$. For the ICL attack Huang et al. (2022a), we fix the number of demonstrations at 4 and run the demonstration selection process using 21 different seeds from a pool of 64 synthetic examples. In the PII Compass attack Nakka et al. (2024), we set the prefix length to 50 tokens and iterate over 64 synthetic prefixes (see Figures 30 and  31). For the SPT attack Kim et al. (2024), we repeat the experiment with 20 task-aware prompt initializations, as shown in Figure 22 in the Appendix.
@@ 8 Ablation Studies
For example, given a real PII pair {Karen Arnold, klarnold@flash.net} in the adversary dataset $\mathcal{D}_{\text{adv}}$ as shown in Figures 24 and 25, we generate synthetic PII pairs in two variations: 1. Altering only the name with email-domain retained (e.g., {"Cameron Thomas", "cthomas@flash.net"}, as shown in Figures 26 and 27 in the Appendix). 2. Altering both the name and the domain with synthetic ones (e.g., {"Cameron Thomas", "cthomas@medresearchinst.org"}, as shown in Figures 28 and 29 in the Appendix).
For synthetic prefixes in the PII-Compass attack (Nakka et al., 2024), we use GPT-3.5 (OpenAI, 2023) to generate email conversation sentences of 50 tokens in length between employees of an energy corporation like Enron, as illustrated in Figures 30 and 31.
The results of PII attacks on these synthetic data experiments are presented in Figures 11 for ICL, PII-Compass, and SPT attacks in three columns, respectively. Overall, our observations are as follows: 1. When both the name and domain are replaced with synthetic data, the extraction rates for both ICL and SPT attacks are notably lower (shown in purple bars) compared to the original performance with real PII pairs (shown in yellow bars). 2. When only the name part is anonymized, the performance of the ICL attack (shown in green bars) remains closer to the original performance with real PII pairs (shown in yellow bars). In contrast, the performance of SPT attacks in this setting shows a significant drop in performance (shown in green bars) from that with original PII pairs (shown in yellow bars) and in fact, the SPT attack, does not even surpass the performance of simple template prompting, as shown in Figure 7(a). 3. With synthetic prefixes generated by GPT (OpenAI, 2023), the performance (shown in purple bars) is substantially lower than the original performance with real prefixes from subjects in $\mathcal{D}_{adv}$, as illustrated in Figure 11(c). Our experiments suggest that for effective PII extraction with PII-Compass, having a prefix that closely resembles the true domain is essential.
Figure 12: Impact of the order of subjects in the demonstration prompt of the ICL attack to extract emails from pretrained GPTJ-6B. We first select $k=\{2,4,6,8,16,32,64\}$ PII pairs from the pool of $M=64$ PII pairs in $\mathcal{D}_{adv}$ using a single seed. Next, we vary the order of the $k$ demonstrations by generating 20 different permutations for each $k$. We visualize the box plot of extraction rates across these 20 different permutations and observe that the ICL attack (Huang et al., 2022a) shows increased sensitivity to demonstration order as the number of demonstrations $k$ increases.
@@
Figure 13: Soft-prompt transferability in Email PII Attacks on pretrained GPTJ-6B. The Y-axis denotes the template structure used for training the soft prompt embeddings. The X-axis shows the four target templates used during the attack stage. To conduct this study, we prepend the trained soft prompt embeddings from different source templates (indicated along the Y-axis) to different target template prompts (indicated along the X-axis) and report the average PII extraction performance over 21 soft-prompt initializations shown in Figure 22.
@@
(a) True-prefix attack
(b) Template attack
Figure 14: True-prefix attack and Template attack on the finetuned GPTJ-6B model to extract Email PII. On the left, we show the performance of the true-prefix attack (Carlini et al., 2021a), and on the right, we present the performance of the template attack (Huang et al., 2022a). Results for the pretrained model are shown in light color, while results for the finetuned model are shown in dark color. Across the board, we observe that PII extraction rates on the finetuned model are significantly higher than those on the pretrained model.
@@
(b) PII-Compass attack
Figure 15: Sensitivity of Hard-Prompt Email PII Attacks on the Finetuned GPTJ-6B Model . Similar to the results on the pretrained model in Figure 7, the ICL attack (Huang et al., 2022a) on the left shows sensitivity to the selection of demonstrations from the available pool of $\mathcal{D}_{adv}$, while the PII Compass attack (Nakka et al., 2024) on the right illustrates the impact of varying true prefixes from other data subjects in $\mathcal{D}_{adv}$.
@@ 9 PII Attacks on Finetuned Model
Figure 16: Sensitivity of SPT Attack (Kim et al., 2024) on the Finetuned GPTJ-6B Model. We examine the variation in PII extraction rates by analyzing the impact of three independent factors. Each factor is varied independently from the base configuration, and the results show that the SPT attack requires careful hyperparameter selection for optimal performance.
@@
Continual PII extraction. We also conduct continual PII extraction on the finetuned model by leveraging successfully extracted PII pairs along with the originally available PII pairs in $\mathcal{D}_{adv}$. We perform this experiment with 5 task-aware initializations (see first 5 in Figure 22 in the Appendix) for each template. From results in Figure 17, we observe that the average extraction rates improve for templates A, B, C, and D from 9.09%, 19.9%, 24.1%, 28.2% at the end of round 1 to 12.1%, 35.8%, 39.5%, 42.1% at the end of round 2. All templates achieve a boost of more than 1.5x, except for template A, which shows greater variance in extraction rates across different initializations. Similar to the findings observed in Figures 7(a)LABEL:fig:icl_attack9(b)14(b)15(a), and 15(b), where different templates exhibit varying extraction rates in different contexts, we also find that template structures, such as A, are less effective in continual PII extraction. Understanding the reasons for this behavior could offer valuable insights into how memorization occurs in LLMs and suggests an avenue for future work focused on optimizing template structures for template structures for PII extraction.
@@ 4 Templates: [A, B, C, D])
Table 5: Evaluating Email PII attacks with higher query budgets on the finetuned GPTJ-6B model. The first four columns outline the threat setting in terms of data access in $\mathcal{D}_{adv}$ and $\mathcal{D}_{eval}$. The fifth column shows the model access type (W.B.: white box, B.B.: black box). We conduct PII attacks by querying the model multiple times, either through simple top-$k$ model sampling or by varying configuration settings within each attack method. Unlike attacks on the pretrained model, even the simple template attack (Huang et al., 2022a) achieves more than 50% accuracy in finetuned settings. Furthermore, similar to earlier results on the pretrained model, we observe that the extraction rate improves by 1.2x-2.8x compared to the best extraction rate observed with a single query.
@@
Figure 17: Continual Email PII extraction on the finetuned GPTJ-6B (Wang & Komatsuzaki, 2021). We report the extraction rates of the SPT attack (Kim et al., 2024) over ten rounds for four templates in a continual learning setting. At the end of each round, successfully extracted PIIs are incorporated to retrain the soft prompt embeddings for the subsequent round. The average extraction rate, along with its range, is plotted for the first five soft-prompt initializations shown in Figure 22.
10 Evaluating Email PII Extraction Attacks on Pythia 6.9B
In this section, we conduct PII extraction attacks on a different model, Pythia 6.9B Biderman et al. (2023), which is also pretrained on the PILE corpus containing the Enron-email dataset (Shetty & Adibi, 2004).
In Tables 6 (a) and (b), we provide the results of repeated queries on the pretrained and finetuned Pythia 6.9B models. Furthermore, Figures 19(a) and (b) show the similar trend of increased email PII extraction rates for the pretrained and finetuned Pythia 6.9B Biderman et al. (2023) models, respectively, in continual settings. Overall, we observe the similar trend of increased extraction rates of upto 4x times in repeated query settings, and as expected, the extraction rates on the finetuned model is higher than the pretrained model. In short, the findings of our underestimation of privacy leakage in single-query settings also generalize to Pythia 6.9B, showcasing the generalization of our results.
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Pretrained model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
13.6% 35.7% (2.6x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
2.6% 10.1% (3.9x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
8.1% 23.4% (2.9x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
7.5% 21.8% (2.9x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
8.8% 28.5% (3.2x) $\uparrow$
(a) Pretrained Pythia 6.9B (Biderman et al., 2023).
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Finetuned model
Pretrained
model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
Pretrained
($K$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$K=256$
38.6% 61.0% (1.2x) $\uparrow$ 35.7%
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$K=256$
14.3% 46.4% (3.2x) $\uparrow$ 10.1%
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$K=440$
23.7% 50.0% (2.1x) $\uparrow$ 23.4%
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$K=164$
25.3% 47.4% (1.8x) $\uparrow$ 21.8%
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$K=768$
25.6% 53.6% (2.1x) $\uparrow$ 28.5%
(b) Finetuned Pythia 6.9B (Biderman et al., 2023)
[…]
(a) Pretrained Pythia 6.9B
(b) Finetuned Pythia 6.9B
[…]
11 Evaluating PII Attacks to Extract Phone Numbers
In this section, we focus on the numerical phone number PII present in the Enron Email dataset (Shetty & Adibi, 2004). To this end, we randomly sample 500 subjects from the 2700 subjects released by the authors in the ICL Attack (Shao et al., 2023). We set aside 64 subjects as the attacker’s knowledge and evaluate the extraction rates on the remaining 436 subjects. For evaluation, we use the exact match metric, where all the numerical digits in the ground truth must match the predicted phone number string. Note that we remove non-numeric characters, such as parentheses and hyphens, before comparing the numbers.
Tables 7 (a) and (b) show the extraction rates with repeated querying on pretrained and finetuned GPT-J-6B (Wang & Komatsuzaki, 2021). Furthermore, Tables 8 (a) and (b) show the results with pretrained and finetuned Pythia 6.9 (Biderman et al., 2023).
Compared to email PII, the extraction rates for phone number PII are lower, which may be partly attributed to the strict evaluation metric of exact match and the more complex nature of phone numbers, which have no direct connection to the subject’s name. In contrast, email PII often includes a user-part that is connected to the subject’s name.
Our experiments with phone number PII also validate our prior findings with email PII with regard to underestimation of privacy leakage in single-query setting and increased extraction rates with repeated querying and in continual settings.
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Pretrained model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
4.1% 11.7% (2.9x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
0.2% 0.5% (2.5x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
1.1% 1.8% (1.6x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
1.6% 4.1% (2.6x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
1.6% 8.2% (5.1x) $\uparrow$
(a) Pretrained GPTJ-6B (Wang & Komatsuzaki, 2021).
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Pretrained model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
29.4% 62.3% (2.1x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
2.5% 11.2% (4.5x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
5.7% 13.3% (2.3x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
6.7% 14.5% (2.2x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
6.2% 20.8% (3.4x) $\uparrow$
(b) Finetuned GPTJ-6B (Wang & Komatsuzaki, 2021).
[…]
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Pretrained model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
4.4% 12.4% (2.8x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
0.2% 0.7% (3.5x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
0.7% 1.8% (2.8x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
1.8% 5.0% (2.8x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
5.7% 15.6% (2.7x) $\uparrow$
(a) Pretrained Pythia 6.9B (Biderman et al., 2023).
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Finetuned model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
12.8% 35.5% (2.8x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
2.5% 11.7% (4.7x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
5.0% 10.7% (2.1x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
5.7% 12.2% (2.1x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
5.7% 17.2% (3.0x) $\uparrow$
(b) Finetuned Pythia 6.9B (Biderman et al., 2023).
[…]
(a) Pretrained GPT-J-6B
(b) Pretrained Pythia 6.9B
(c) Finetuned GPTJ-6B
(d) Finetuned Pythia 6.9B
[…]
12 PII Attacks on LLaMa-7B with PII-Scrubbing Defense
To defend against PII attacks, PII-scrubbing— a technique that detects and removes PII entities— is commonly used to enhance data privacy in pretraining scenarios, valued for its efficiency and reasonable performance. In this section, we evaluate PII attacks on LLaMa-7B (Touvron et al., 2023) finetuned on the PII-scrubbed Enron-email dataset using the Flair toolbox (Akbik et al., 2019), and benchmark it against LLaMa-7B finetuned without scrubbing. We use the publicly provided finetuned models released by the authors of LLM-PBE (Li et al., 2024b).
We present the results of our evaluation in repeated query settings in Table 9, and in continual settings in Figure 20. As expected, the scrubbed model has lower extraction rates compared to the undefended model. However, our findings show that the single-query attack severely underestimates the privacy leakage, even for the scrubbing defense. In fact, the relative boost in extraction rates is much higher for the scrubbing-based finetuned model. For example, the improvement in extraction rates with repeated querying is 4.1x and 5x for the scrubbed model, compared to 1.7x and 2.8x for the undefended model on SPT and Template attacks, respectively. Finally, the extraction rates in continual settings for the scrubbed model also increases over the rounds, albeit the absolute extraction rates are lower than the undefended model as shown in Figure 20.
Attacker’s Knowledge
in $\bm{D_{adv}}$
Attacker’s Knowledge
of query $q$ data subject
in $\bm{D_{eval}}$
Finetuned model
True-prefix
$\bm{\{r_{j}\}_{j=1}^{M}}$
PII pairs
$\bm{\{s_{j},p_{j}\}_{j=1}^{M}}$
True-prefix
$\bm{r_{q}}$
Subject name
$\bm{s_{q}}$
Data
Scrubbing
Model
access
PII Attack
Model
Sampling
Number of
Queries
Accuracy
(1 query,
best case)
Accuracy
($k$-queries)
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
37.7% 60.4% (1.6x) $\uparrow$
$\circ$
$\circ$
$\bullet$
$\circ$
B.B
True-prefix
 (Carlini et al., 2021a)
$k=256$
15.2% 34.4% (2.3x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
10.7% 30.2 % (2.8x) $\uparrow$
$\circ$
$\circ$
$\circ$
$\bullet$
B.B
Template
 (Huang et al., 2022a)
$k=256$
3.5% 17.2% (5.0x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
13.6% 40.6% (2.9x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
B.B
ICL
 (Huang et al., 2022a)
$k=440$
4.9% 4.8% (2.88x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
27.3% 47.4% (1.7x) $\uparrow$
$\circ$
$\bullet$
$\circ$
$\bullet$
W.B
SPT
 (Kim et al., 2024)
$k=164$
5.5% 22.4% (4.1x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
13.6% 44.1% (3.2x) $\uparrow$
$\bullet$
$\circ$
$\circ$
$\bullet$
B.B
PII Compass
 (Nakka et al., 2024)
$k=768$
5.8% 31.2% (5.4x) $\uparrow$
[…]